Determine the capitalized cost of a nuclear power plant which requires P20,000,000 for original
construction; P350,000 at the end of every year for the first 5 years and then P200,000 each year
thereafter for operating expenses, and P1,000,000 every 3 years for replacement of equipment with
interest at 16% per annum.
1
Expert's answer
2021-09-26T11:56:27-0400

Capitalized cost

"= 5,000,000 + 100,000(\frac{P}{A} 12 \%, 6) + (\frac{120,000 }{ 0.12}) (\frac{P}{F}, 12 \%, 6) + \frac{[500,000(\frac{A}{F}, 12\%, 5)]}{0.12}\\

 = 5,000,000 + 100,000(4.111) + 1,000,000(0.5066) + \frac{[500,000(0.1574)]}{ 0.12}\\

 = 5,000,000 + 411,100 + 506,600 + 655,833.33\\

 = \$6,573,533.33"